The pH of a 0.0147M KCN solution is

A 8.48
B 3.3
C 5.52
D

10.7
Answer #1

Ken - kt & EN- (N- & 120 e non 0.0147 - 4x to - x 0.0147 - x AAAAAYYY77 x 3 Kb = (0.0147-2) ( 1,6X600) (0.0147) x2 x = 4.84 X